ABC Communications recently completed a 5-for-3 stock split. Prior to the split, its stock price was $70 per share. The firm's total market value increased by 20% as a result of the split. What was the price of the company’s stock following the stock split?

a.

$46.2

b.

$67.2

c.

$61.6

d.

$50.4
Answer #1

Stock price before split=$70

Each of 3 shares became 5 shared after split...

Stock price after split will be=3(70)÷5

=210÷5

=42...

Increase in market value after split =20℅

Therefore market price of stock after split=42+20℅=50.4 per share...

Answer will be

d)50.4 per share...